5. At 1000 degrees * C , water can be decomposed into hydrogen and oxygen gas : 2H 2 O(g) longrightarrow2H 2 (g)+O 2 (g) Calculate the hydrogen and oxygen concentrations at equilibrium if the initial concentration of the water is 0.500 M and Keq = 7.3 * 10 ^ - 18


1
Expert's answer
2021-02-28T06:19:58-0500

Keq=[H2]^2 x [O2]/[H2O]^2

If [O2]=x, [H2]=2x.

7.3 *10-18=x * 2x^2/0

5.

2x^3=3.65*10-18

x= 1,22204619e-6 M.

1,22204619e-6 M of O2.

1,22204619e-6 x 2 = 2,44409238e-6 M of H2.